# Large Language Models as Carriers of Hidden Messages

### v2

@@ VII Conclusion
To the best of our knowledge, this is the first publication that proposes a paradigm for extracting LLM fingerprint without the need for infeasible trigger guessing. Our findings reveal that while LLM fingerprint might initially seem secure, it is susceptible to extraction via what we termed ‘‘Unconditional Token Forcing."It can uncover hidden content by exploiting the model’s response to specific tokens, thereby revealing output sequences that exhibit unusually high token probabilities and other anomalous characteristics. Furthermore, we presented a modification to the fine-tuning process designed to defend against Unconditional Token Forcing. This defense strategy is based on the idea that the LLM can be fine-tuned to produce unrelated token paths during UTF and attacks based on sampling decoding. To the best of our knowledge, there is no extraction attack method that can reveal text hidden in LLM with UTFC method.
@@
The methods presented in this publication are a double-edged sword. On one hand, the proposed Unconditional Token Forcing Confusion (UTFC) technique can enhance the robustness of LLM fingerprinting. On the other hand, the same method can be exploited for LLM steganography, enabling covert communication channels that could be used for malign purposes. We believe it is better to openly publish these methods and highlight the associated security concerns so that the community can develop solutions to address those concerns.

### v3

@@ VII Conclusion
To the best of our knowledge, this is the first publication that proposes a paradigm for extracting LLM fingerprint without the need for infeasible trigger guessing. Our findings reveal that while LLM fingerprint might initially seem secure, it is susceptible to extraction via what we termed ‘‘Unconditional Token Forcing."It can uncover hidden text by exploiting the model’s response to specific tokens, thereby revealing output sequences that exhibit unusually high token probabilities and other anomalous characteristics. Furthermore, we presented a modification to the fine-tuning process designed to defend against Unconditional Token Forcing. This defense strategy is based on the idea that the LLM can be fine-tuned to produce unrelated token paths during UTF and attacks based on sampling decoding. To the best of our knowledge, there is no extraction attack method that can reveal text hidden in LLM with UTFC method.
@@
The methods presented in this publication are a double-edged sword. On one hand, the proposed Unconditional Token Forcing Confusion (UTFC) technique can enhance the robustness of LLM fingerprinting. On the other hand, the same method can be exploited for LLM steganography, enabling covert communication channels that could be used for malign purposes. We believe it is better to openly publish these methods and highlight the associated security concerns so that the community can develop solutions to address them.